OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5/9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 8 degrees Celsius is.

OpenStudy (anonymous):

(a)\[ C = \dfrac {5}{9} (F - 32) \rightarrow F - 32 = \dfrac {9}{5}C \rightarrow \boxed {F = \dfrac {9}{5}C + 32}. \](b) \[ F = \dfrac {9}{5} \cdot 8 + 32 = \dfrac {72}{5} + 32 = 14.4 + 32 = \boxed {46.4}. \](In degrees Fahrenheit.)

OpenStudy (anonymous):

The first thing I did is divide the fraction (5/9) to both sides C=(5/9)*(F-32) (5/9) (5/9) ------------------- C*(9/5) = F-32 Then I added 32 to the problem: C*(9/5) + 32 = F For part 2 I just plugged in any value for C to get my value of F. [(2)*9/5] + 32 = F 3.6+ 32 = F 35.6 = F Is this right